OpenStudy (anonymous):

How do I reference an {equation} with the same equation number later in the document? I have: \begin{equations} \begin{align} y = mx + b \end{align} \end{equations} And it shows that it's equation (1) in my document. How do I repeat that same equation later in my document after other equations with the same equation number (1)?

OpenStudy (poopsiedoodle):

maybe you could copy the latex code from the equation and pate it in there later

OpenStudy (anonymous):

I did but then it shows the equation number as something else like (5) or something

OpenStudy (kinggeorge):

``` \begin{equation} \label{eq1} \begin{align} y = mx + b \end{align} \end{equations} ``` By adding the \label{eq1}, you can later type \ref{eq1} and it should reference your equation

OpenStudy (kinggeorge):

Those question marks should be `\ref{eq1}`.